OpenStudy (ittybitty14):

PLEASE HELP!!!!! FAN & MEDAL! :))) If r = 9 units and x = 8 units, then what is the volume of the cylinder shown above? A. 2,592 cubic units B. 306 cubic units C. 72 cubic units D. 648 cubic units

OpenStudy (ittybitty14):

@rebeccaxhawaii @Zelda @studygirl13

OpenStudy (anonymous):

Uh.. picture of cylinder please?

OpenStudy (anonymous):

were is the cyclinder

OpenStudy (anonymous):

@IttyBitty14 Although the volume of the cylinder is given by: \[pi*r^2*h\] In this case h = x. So \[pi*r^2*x = 3.14*r*r*x = V\]

OpenStudy (anonymous):

ok multiple x time r

OpenStudy (anonymous):

and 3.14

OpenStudy (anonymous):

@studygirl13 And another r. ;)

OpenStudy (anonymous):

V=πr2h this is the volume of a cyclinder

OpenStudy (anonymous):

so 3.149^28

OpenStudy (anonymous):

so what is 9 times 9

OpenStudy (anonymous):

@IttyBitty14

OpenStudy (ittybitty14):

81 sorry my internet is slow right now

OpenStudy (anonymous):

ok so 3.14 time 81

OpenStudy (ittybitty14):

254.34

OpenStudy (ittybitty14):

@studygirl13

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

ok so now 214.34 times8

OpenStudy (anonymous):

are u sure u gave me the right numbers

OpenStudy (ittybitty14):

1714.72

OpenStudy (anonymous):

because that is how you solve it but its not one of the answers

OpenStudy (anonymous):

oh oops i made a mistake

OpenStudy (ittybitty14):

254.34 times 8

OpenStudy (anonymous):

V=πr2h oops because you are supposed to find the hieght

OpenStudy (ittybitty14):

okay

OpenStudy (anonymous):

3.14time 81=214.34 but then we have to find the hieght

OpenStudy (ittybitty14):

okay! :)

OpenStudy (ittybitty14):

2034.72

OpenStudy (anonymous):

but it's not one of your answers

OpenStudy (anonymous):

or is this 2034.72 just the hight

OpenStudy (anonymous):

sorry but i gtg